two planes, which are 3400 miles apart, fly toward each other .their speeds differ by 50mph.if they pass each other in 4 hours, what is the speed of each?
Answer by macston(5194)

.
3400 mi/4hrs = 850 mph
.
The combined speeds of both planes is 850 mph.
.
S=speed of slower plane; F=speed of faster plane=S+50 mph
.
S+F=850mph
S+(S+50mph)=850mph
2S=800mph
S=400mph
F=S+50mph=400mph+50mph=450mph
.
ANSWER: The slower plane is traveling 400 mph and the
faster plane is traveling 450 mph.
.